OpenStudy (anonymous):

3^2x-3^(x+2)+8=0 Solve for x.

OpenStudy (shubhamsrg):

let 3^x = t

OpenStudy (anonymous):

\[3^{2x}-3^{x+2}+8=0\] is it that??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

follow @shubhamsrg

OpenStudy (anonymous):

what do you do to 3^(x+2)

OpenStudy (anonymous):

3^(x+2)=(3^x)*(3^2)

OpenStudy (anonymous):

it's a quadratic equation.

OpenStudy (anonymous):

you both deserve a medal. thanks a lot!

OpenStudy (anonymous):

welcome dude :)